OpenStudy (anonymous):

Solve K = (b – 2a) for b. show all work.

OpenStudy (candyme):

You would move the 2a to the other side so it would be k+2a=b

OpenStudy (anonymous):

To be honest the best I have is this: K=b-2a Move the 2a over The answer is 2a+k=b

OpenStudy (candyme):

Because to move it you subtract a negative 2a from both sides, which is basically adding a positive to both sides

OpenStudy (candyme):

You have it right! :)

OpenStudy (anonymous):

thanks a bunch i don't have learning material for summer school so this is helping me a lot you guys are awesome!